A rectangle is 4 meters longer than it is wide. The area of the rectangle is 252 square meters. Find the length and width of the rectangle.
length = ________ m width = ________ m

Answer by nerdybill(7384)

A rectangle is 4 meters longer than it is wide. The area of the rectangle is 252 square meters. Find the length and width of the rectangle.
.
Let w = width of rectangle
then
w+4 = length of rectangle
.
w(w+4) = 252
w^2+4w = 252
w^2+4w-252 = 0
(w-14)(w+18) = 0
w = {14, -18}
Throw out the negative solution leaving:
w = 14 meters (width)
.
w+4 =14+4 = 18 meters (length)